hey guys I am really struggling with something at the moment, ill show you.
it has to with the - + signs.
if I have 2-5x=3x+2
am I reading the equation as -5x -3x = 2-2? or am I reading it as 5x-3x= 2-2.
super confused at the moment would be great if you could help me out
Cheers jeremy

Answer by mananth(16946)

for functional knowledge
whenever you move the term from right side of equation to left side or vice versa
the sign changes
3x-2 =x-4
3x-x=-4+2




-5x -3x = 2-2 this is right